# ExMAG: Learning of Maximally Ancestral Graphs

### v2

@@ Abstract
As one transitions from statistical to causal learning, one is seeking the most appropriate causal model. Dynamic Bayesian networks are a popular model, where a weighted directed acyclic graph represents the causal relationships. Stochastic processes are represented by its vertices, and weighted oriented edges suggest the strength of the causal relationships. When there are confounders, one would like to utilize both oriented edges (when the direction of causality is clear) and edges that are not oriented (when there is a confounder or not a relationship), yielding mixed graphs. A little-studied extension of acyclicity to this mixed-graph setting is known as maximally ancestral graphs with consideration of confounders.
We propose a score-based learning algorithm for learning maximally ancestral graphs. A mixed-integer quadratic program is formulated, and an algorithmic approach is proposed, in which the pre-generation of exponentially many constraints is avoided by generating only violated constraints in the so-called branch-and-cut (“lazy constraint”) method. Comparing the novel approach to the state-of-the-art, we show that the proposed approach turns out to produce more accurate results when applied to small and medium-sized synthetic instances containing up to 25 variables.
@@ 1 Introduction
As one transitions from statistical to causal learning [47], one is seeking the most appropriate causal model. Dynamic Bayesian networks (DBN) [17, 35] are a popular model, where a weighted directed acyclic graph represents the causal relationships. Stochastic processes are represented by its vertices, and weighted oriented edges suggest the strength of the causal relationships. The key challenge in learning DBNs is confounding.
To illustrate the challenge of confounding, let us consider Simpson’s paradox. Simpson’s paradox shows that without considering confounding factors in statistical analysis [34], the direction of causality can be mis-estimated completely. An textbook example [34] comes from the Berkeley graduate admissions [3]. The data show that women find it harder to get admitted to Berkeley graduate schools. Nevertheless, this is because women tend to apply to departments that have lower admission rates. In this example, the choice of the graduate school is the confounder, impacting the probability of admission. Confounding is prevalent throughout high-dimensional statistics [32, 21, e.g.], such as in biomedical sciences.
@@
In statistical theory, [8] study confounding in detail, and many subsequent works develop further methods. [31] show that leveraging the dominant eigenstructure of time series may improve performance of estimation. Anchor regression, for instance, bridges the gap between causality and robustness by addressing heterogeneity in data [44]. Other significant contributions include spectral deconfounding, a technique designed to mitigate the effects of hidden confounders in high-dimensional settings [7]. This approach provides a framework for robust predictions in the presence of shifts in data distributions. Similarly, the invariance principle has emerged as a cornerstone of causal inference, linking causal structure to robust statistical models [6]. Furthermore, the concept of doubly robust inference offers an alternative framework for addressing hidden confounding factors, combining model robustness with efficiency in high-dimensional scenarios [22]. Together, these developments represent a significant step forward in understanding and addressing the challenges posed by complex causal systems.
@@
While there is a long history of the study of confounding, as suggested above, the extensions of DBNs to allow for confounding are rather more recent. Instead of estimating a Directed Acyclic Graph (DAG), one could estimate a Maximal Ancestral Graph (MAG), cf. [41]. These allow for both direct and indirect relationships among variables modelled as directed and undirected edges, even in the presence of confounding factors. In particular, MAGs can represent feedback loops and bidirectional relationships that DAG-based models, such as DBNs, cannot. This makes MAGs a more powerful tool for capturing the complex dynamics of real-world causal systems.
@@
There are only a few studies of MAG estimation [10, 38, 12, 24, 27, 26, 15]. [40, 49] are applicable to the both discrete and nonparametric cases, which extend DAG to MAG, ADMG or Bow-Free Acyclic Path Diagrams, see more definition in 2. Factorization in MAG is not directly decomposable into individual variables and their parent sets, as in DAGs, but must instead consider components connected by bidirected paths (termed districts or c-components), cf. [40], although [12] proposed to use Markov equivalence classes (MEC) instead. In 2021, [10] introduced a first mixed-integer programming (MIP) formulation, but the number of variables scales with the number of c-components, i. e., exponentially with the number of vertices in the worst case. Such formulations are also known as extended formulations [13]. The same year, [38] explored a constraint-based approach for MAG discovery, leveraging conditional independence testing. Additionally, [53] addresses exogenous covariates in causal formulation that helps explain the heterogeneity in both sampling and causal mechanisms. The dissertation of [24] presented an extension of the imsets of [48] from directed acyclic graphs (DAGs) to towards MAGs [27], which allows for the use of the methods of Studený, and a score-based heuristic [26]. More recently, the paper by [15] enhanced the scalability of methods of [10] by utilizing linear programming (LP) relaxations instead of solving the MIP.
@@
Our approach proposes the first compact formulation of MAG estimation within Mixed-Integer Nonlinear Programming (MINLP), in contrast with the so-called extended formulations of [10, 15]. While both the extended and compact formulations ensure that confounding factors are properly accounted for and the true underlying data-generating process is better represented by the model, our implementation scales further, from 4-5 stochastic processes in the extended formulation to 25 or more stochastic processes with the proposed compact formulation.
@@ 1.1 Motivating Example
Let us revisit the Berkeley graduate admission paradox example of page 1. As in most paradoxes, there is no violation of logic in Simpson’s paradox, just a violation of of intuition. The poor intuition being violated in this case is that a positive association in the entire population should also hold within each department. Overall, females in these data did have a harder time getting admitted to graduate school. But that arose, because female applicants chose the departments that were the most difficult to gain admission to for anyone, male or female. In this example, gender influences the choice of department, and the department influences the chance of admission. Controlling for department reveals a more plausible direct causal influence of gender. If one represents gender as $G$, department as $D$, acceptance as $A$, and academic ability as $U$, the potential confounder, the DAG form is
[…]
Our method, ExMAG, is able to reveal the confounders in this Berkeley graduate admission example, as illustrated in the notebook in supplementary materials and Figure 1.
Figure 1: Performance of ExDBN of [45] (top) and ExMAG (bottom) on the Berkeley graduate admission example. While the DBN identified by ExDBN suggests a causal relationship between gender and admission, ExMAG correctly identifies the confounding.
@@ 2 Graphs and Properties
Maximal ancestral graphs (MAGs), first introduced by [41], provide a framework for modeling distributions through conditional independence (CI) relations. Compared with directed acyclic graphs (DAGs), MAGs allow for latent confounders, accommodating data that arises from distributions with more complex independence structures and revealing hidden states in the graphs. While DAGs allow for the efficient computation of maximum likelihood estimates (MLEs) and scoring (e.g., via BIC), these properties are challenging to extend to MAG due to their structural and computational complexity [25].
Directed Acyclic Graph
A DAG is a directed graph $G=(V,E)$ such that there are no directed cycles. That is, there is no sequence of distinct vertices $v_{1},v_{2},\dots,v_{k}\in V$ such that $(v_{i},v_{i+1})\in E$ for all $11\leq i\leq k-$ 1 and $(v_{k},v_{1})\in E$.
Bow-Free Acyclic Path Diagrams
Let $G$ be a bow-free ADMG (Acyclic Directed Mixed Graph). The graph $G$ is bow-free if and only if there does not exist a directed and bidirected edge between a pair of adjacent vertices. More formally, for any $v_{i},v_{j}\in V$, there cannot exist both $v_{i}\to v_{j}$ (or $v_{i}\leftarrow v_{j}$) and $v_{i}\leftrightarrow v_{j}$ simultaneously, as described by [50] .
@@
$\text{pa}_{\mathcal{G}}(v)=\{w:w\to v\text{ in }\mathcal{G}\},$
$\text{sp}_{\mathcal{G}}(v)=\{w:w\leftrightarrow v\text{ in }\mathcal{G}\}$
$\text{an}_{\mathcal{G}}(v)=\{w:w\to\cdots\to v\text{ in }\mathcal{G}\text{ or }w=v\},$
$\text{dis}_{\mathcal{G}}(v)=\{w:w\leftrightarrow\cdots\leftrightarrow v\text{ in }\mathcal{G}\text{ or }w=v\}.$
Given a directed mixed graph $\mathcal{G}$, the districts define a set of equivalence classes of nodes in $\mathcal{G}$. The district for node $v$ is defined as the connected component of $v$ in the subgraph of $\mathcal{G}$ induced by all bidirected edges. As in a DAG, a mixed graph $\mathcal{G}$ is acyclic if it contains no directed cycles in $\mathcal{E}_{d}$, i.e., an acyclic directed mixed graph (ADMG) [24].
@@
$\text{If }v\neq w\text{ and }w\in\text{an}_{\mathcal{G}}(v)\cup\text{sp}_{\mathcal{G}}(v),\text{ then }v\notin\text{an}_{\mathcal{G}}(w),$
which is written as, $\mathcal{G}$ is an ancestral ADMG if it contains no directed cycles ($v\to u\to\ldots\to w\to v$) or almost directed cycles [10, 24].
An almost directed cycle is of the form $v\to u\to\ldots\to w\leftrightarrow v$; in other words, $\{v,w\}\in\mathcal{E}_{w}$ is a bidirected edge, and $v\in\text{an}_{\mathcal{G}}(w)$ [10].
@@
An inducing path from a variable $X$ to a variable $Y$ in a directed graph $G=(V,E)$ is a path $P=(X=v_{0},v_{1},\dots,v_{n}=Y)$ such that for all intermediate nodes $v_{i}$ (where $11\leq i\leq n-$ 1), $v_{i}\in Z$, where $Z$ is the conditioning set. If the path is blocked by conditioning on $Z$, then it is an inducing path.
A node j on a non-overlapping path is called a collider if contains a non-overlapping subpath (i; j; k) with two arrowheads into j. In mathematical form, collider is represented as
$\text{collider}_{G}(u)=\left\{u:\exists w,v:\begin{array}[]{c}w\to u\leftarrow v\vee\\ w\leftrightarrow u\leftarrow v\vee\\ w\to u\leftrightarrow v\vee\\ w\leftrightarrow u\leftrightarrow v\end{array}\right\}.$
@@
For an ADMG $\mathcal{G}$, given a subset $W\subseteq\mathcal{V}$, the induced subgraph $\mathcal{G}_{W}$ is defined as the graph with vertex set $W$ and edges in $\mathcal{G}$ whose endpoints are both in $W$. Also for the district of a vertex $v$ in an induced subgraph $\mathcal{G}_{W}$, we may denote it by $\text{dis}_{W}(v)=\text{dis}_{G_{W}}(v)$.
Graphs are associated with conditional independence relations via a separation criterion; in the case of ADMGs, we use m-separation. Graphs encode conditional independence relations through separation criteria. For ADMGs, m-separation is used, which generalizes d-separation for DAGs to account for bidirected edges. A path between two vertices $a$ and $b$ is m-connecting given a conditioning set $C\subseteq\mathcal{V}$ if:
  1. 1.
    $a$ and $b$ are endpoints of the path;
  2. 2.
    Every noncollider on the path is not in $C$;
  3. 3.
    Every collider is in $\text{an}_{G}(C)$.
Graphs that encode the same set of CI relations are said to belong to the same MEC. This equivalence class structure is critical for reducing redundancy during graph exploration and scoring.
@@
We refer to [24] for multiple examples.
3 Brief Introduction To Mixed Integer Quadratic Programming
$\underset{x\in\mathbb{R}^{n}}{\min}\quad$ $x^{T}Qx+q^{T}x,$ (1)
s.t. $x^{T}Q_{i}x+q_{i}^{T}x\leq a_{i},$ (2)
$Ax\leq b,$ (3)
$x\in F$ (4)
where $Q,Q_{i}\in\mathbb{R}^{n,n}$, $q,q_{i}\in\mathbb{R}^{n}$, $A\in\mathbb{R}^{m,n}$, $a\in\mathbb{R}^{k}$, $b\in\mathbb{R}^{m}$, $F$ is a product of the form
$F=\underbrace{\mathbb{R}\times\ldots\times\mathbb{R}}_{n-r\text{ times}}\times\underbrace{\mathbb{N}\times\ldots\times\mathbb{N}}_{r\text{ times}}$ (5)
and $m,n,k,r\in\mathbb{N}$. In the mathematical form, (1) is often called the cost or loss function, (2) represents the quadratic constraints, (3) is the linear constraints, and $F$ is the set that enforces the integrality constraints for the $r$ components of the decision variable $x$.
Due to the exhaustive nature of the algorithms mentioned in the previous paragraph, global convergence is guaranteed [2]. Furthermore, convergence to the global solution may be tracked, and the error estimated by computing the dual problem of (14). The dual of the problem is then used to compute the so-called MIP GAP as follows
@@ 4.1 Connecting to Causality
In this perspective, we present the formulation of the Mixed Integer Quadratic Program (MIQP) used to infer the causal structure, with a new binary matrix $B=[b_{j,k}]\in\{0,1\}^{d\times d}$ introduced to account for relationships explained by confounding factors, alongside the weight matrix $W$ and binary adjacency matrix $E$ adopted from the ExDAG model [46]. Weight matrix $W$ represents the model weights. Whenever an entry in $W_{j,k}$ is non-zero, the respective value in either $E_{j,k}$ (directed edge) or $B_{j,k}$ is nonzero (bidirected edge). At the same time, we extend the existing formulation by introducing an additional binary input parameter $f_{j,k}$ for each pair of variables $(j,k)$, where $j\neq k$, indicating that there is no direct causal relationship between variables $j$ and $k$, but $j$ and $k$ might have a common cofactor. This follows from the meaning of the edges in a MAG $\rightarrow$ edge implies a direct causal relationship but does not rule out a possible latent confounding, $\leftrightarrow$ means no direct causal relationship. Formally, we define the Directed Edge Matrix $E$ as
$e_{j,k}=\begin{cases}1,&\text{if $j\to k$},\\ 0,&\text{otherwise},\end{cases}$ (10)
and the Bidirected Edge Matrix $B$ as
$b_{j,k}=\begin{cases}1,&\text{if $j\leftrightarrow k$},\\ 0,&\text{otherwise}.\end{cases}$ (11)
The new input matrix $F$, is by the definition,
$f_{j,k}=\begin{cases}1,&\mbox{if }j\not\to k\quad\mbox{and}\quad k\not\to j,\\ 0,&\mbox{otherwise}.\end{cases}$ (12)
Based on these assumptions, one can substitute algebraic equalities 9 into 8, and the work of [46] is extended as explained in the next section.

### v3

@@ Abstract
In mixed graphs, there are both directed and undirected edges. An extension of acyclicity to this mixed-graph setting is known as maximally ancestral graphs. This extension is of considerable interest in causal learning in the presence of confounders. There, directed edges represent a clear direction of causality, while undirected edges represent confounding.
We propose a score-based branch-and-cut algorithm for learning maximally ancestral graphs. The algorithm produces more accurate results than state-of-the-art methods, while being faster to run on small and medium-sized synthetic instances.
@@ 1 Introduction
As one transitions from statistical to causal learning [50], one is seeking the most appropriate causal model. Dynamic Bayesian networks (DBN) [19, 38] are a popular model, where a weighted directed acyclic graph represents the causal relationships. Stochastic processes are represented by their vertices, and weighted, oriented edges suggest the strength of the causal relationships. The key challenge in learning DBNs is confounding.
To illustrate the challenge of confounding, let us consider Simpson’s paradox. Simpson’s paradox shows that without considering confounding factors in statistical analysis [36], the direction of causality can be mis-estimated completely. An textbook example [36] comes from the Berkeley graduate admissions [4]. The data show that women find it harder to get admitted to Berkeley graduate schools. Nevertheless, this is because women tend to apply to departments that have lower admission rates. In this example, the choice of the graduate school is the confounder, impacting the probability of admission. Confounding is prevalent throughout high-dimensional statistics [34, 23], such as in biomedical sciences.
@@
In statistical theory, work [9] studies confounding in detail, and many subsequent works develop further methods. [33] shows that leveraging the dominant eigenstructure of time series may improve performance of estimation. Anchor regression, for instance, bridges the gap between causality and robustness by addressing heterogeneity in data [47]. Other significant contributions include spectral deconfounding, a technique designed to mitigate the effects of hidden confounders in high-dimensional settings [8]. This approach provides a framework for robust predictions in the presence of shifts in data distributions. Similarly, the invariance principle has emerged as a cornerstone of causal inference, linking causal structure to robust statistical models [7]. Furthermore, the concept of doubly robust inference offers an alternative framework for addressing hidden confounding factors, combining model robustness with efficiency in high-dimensional scenarios [24]. Together, these developments represent a significant step forward in understanding and addressing the challenges posed by complex causal systems.
@@
While there is a long history of the study of confounding, as suggested above, the extensions of DBNs to allow for confounding are rather more recent. Instead of estimating a Directed Acyclic Graph (DAG), one could estimate a Maximal Ancestral Graph (MAG), cf. [44]. MAGs allow for both direct and indirect relationships among variables modelled as directed and undirected edges, even in the presence of confounding factors. In particular, MAGs can represent feedback loops and bidirectional relationships that DAG-based models, such as DBNs, cannot. This makes MAGs a more powerful tool for capturing the complex dynamics of real-world causal systems.
@@
There are only a few studies of MAG estimation [11, 41, 13, 26, 29, 28, 17]. [43, 53] are applicable to both discrete and nonparametric cases, which extend DAG to MAG or ADMG diagrams, see more definitions in Sect. 2. Factorization in MAG is not directly decomposable into individual variables and their parent sets, as in DAGs, but must instead consider components connected by bidirected paths (termed districts or c-components), cf. [43], although [13] proposed to use Markov equivalence classes (MEC) instead. In 2021, [11] introduced a first mixed-integer programming (MIP) formulation, but the number of variables scales with the number of c-components, i. e., exponentially with the number of vertices in the worst case. Such formulations are also known as extended formulations [14]. The same year, [41] explored a constraint-based approach for MAG discovery, leveraging conditional independence testing. Additionally, [56] addresses exogenous covariates in causal formulation that helps explain the heterogeneity in both sampling and causal mechanisms. Dissertation [26] presented an extension of the imsets of [52] from directed acyclic graphs (DAGs) to towards MAGs [29], which allows for the use of the methods of Studený, and a score-based heuristic [28]. More recently, paper [17] enhanced the scalability of methods of [11] by utilizing linear programming (LP) relaxations instead of solving the MIP.
@@
Our approach proposes a formulation of MAG estimation within Mixed-Integer Nonlinear Programming (MINLP) in a dimension polynomial in the number of vertices, in contrast with the so-called extended formulations of [11, 17], where the dimension is exponential in the number of vertices. While both the extended formulation of [11] and ours ensure that confounding factors are properly accounted for and the true underlying data-generating process is better represented by the model, our implementation scales further, from 4-5 stochastic processes in the extended formulation of [11] to 25 or more stochastic processes with the proposed compact formulation.
AbilityDepartmentAdmitGender
(a) Cofounding graph
[…]
(b) Result of ExDBN
[…]
(c) Result of ExMAG
Figure 1: Ground truth with the confounder of Department on the Berkeley graduate admission example (left, 1(a)), a dynamic Bayesian network trained on the data (center, 1(b)), and ExMAG output (right, 1(c)). While the dynamic Bayesian network suggests a causal relationship between gender and admission, ExMAG correctly identifies the confounding. See the supplementary material for details.
@@ 1.1 Motivating Example
Let us revisit the Berkeley graduate admission paradox example from the first page. As in most paradoxes, there is no violation of logic in Simpson’s paradox, just a violation of of intuition. The poor intuition being violated in this case is that a positive association in the entire population should also hold within each department. Overall, females in these data did have a harder time getting admitted to graduate school. But that arose, because female applicants chose the departments that were the most difficult to gain admission to for anyone, male or female. In this example, gender influences the choice of department, and the department influences the chance of admission. Controlling for department reveals a more plausible direct causal influence of gender, as illustrated in the DAG in Fig. 1(a). Our method, ExMAG, is able to reveal the confounders in this Berkeley graduate admission example, as illustrated in the notebook in supplementary materials and Figure 1.
@@ 2 Graphs and Properties
A DAG is a directed graph $G=(V,E)$ such that there are no directed cycles. That is, there is no sequence of distinct vertices $v_{1},v_{2},\dots,v_{k}\in V$ such that $(v_{i},v_{i+1})\in E$ for all $11\leq i\leq k-$ 1 and $(v_{k},v_{1})\in E$. Maximal ancestral graphs (MAGs), first introduced by [44], provide a framework for modeling distributions through conditional independence (CI) relations. Compared with directed acyclic graphs (DAGs), MAGs allow for latent confounders, accommodating data that arise from distributions with more complex independence structures and revealing hidden states in the graphs. While DAGs allow for the efficient computation of maximum likelihood estimates (MLEs) and scoring (e.g., via BIC), these properties are challenging to extend to MAG due to their structural and computational complexity [27].
@@
$\text{pa}_{\mathcal{G}}(v)$ =$=\{w:w\to v\text{ in }\mathcal{G}\},$ $\text{an}_{\mathcal{G}}(v)$ =$=\{w:w\to\cdots\to v\text{ in }\mathcal{G}\text{ or }w=v\},$
$\text{sp}_{\mathcal{G}}(v)$ =$=\{w:w\leftrightarrow v\text{ in }\mathcal{G}\},$ $\text{dis}_{\mathcal{G}}(v)$ =$=\{w:w\leftrightarrow\cdots\leftrightarrow v\text{ in }\mathcal{G}\text{ or }w=v\}.$
Given a directed mixed graph $\mathcal{G}$, the districts define a set of equivalence classes of nodes in $\mathcal{G}$. The district for node $v$ is defined as the connected component of $v$ in the subgraph of $\mathcal{G}$ induced by all bidirected edges. As in a DAG, a mixed graph $\mathcal{G}$ is acyclic if it contains no directed cycles in $\mathcal{E}$, i.e., an acyclic directed mixed graph (ADMG) [26].
@@
$\text{If }v\neq w\text{ and }v\in\text{an}_{\mathcal{G}}(w)\cup\text{sp}_{\mathcal{G}}(w),\text{ then }w\notin\text{an}_{\mathcal{G}}(v),$
which is written as, $\mathcal{G}$ is an ancestral ADMG if it contains no directed cycles ($v\to u\to\ldots\to w\to v$) or almost directed cycles [11, 26]. In an ADMG, an almost directed cycle is of the form $v\to u\to\ldots\to w\leftrightarrow v$; in other words, $\{v,w\}\in\mathcal{U}$ is a bidirected edge, and $v\in\text{an}_{\mathcal{G}}(w)$ [11].
@@
An inducing path from variable $X$ to variable $Y$ in a directed graph $G=(V,E)$ is a path $P=(X=v_{0},v_{1},\dots,v_{n}=Y)$ such that for all intermediate nodes $v_{i}$ (where $11\leq i\leq n-$ 1), $v_{i}\in Z$, where $Z$ is the conditioning set. If the path is blocked by conditioning on $Z$, then it is an inducing path.
A node $u$ on a non-overlapping path is called a collider if contains a non-overlapping subpath $(w,u,v)$ with two arrowheads into $u$. In mathematical form, a collider is represented as
$\text{collider}_{G}(u)=\left\{u:\exists w,v:\begin{array}[]{c}w\to u\leftarrow v\vee w\leftrightarrow u\leftarrow v\vee w\to u\leftrightarrow v\vee w\leftrightarrow u\leftrightarrow v\end{array}\right\}.$
@@
Graphs encode conditional independence via separation criteria. For acyclic directed mixed graphs (ADMGs), m-separation generalizes d-separation to handle bidirected edges. A path between nodes $a$ and $b$ is m-connecting given a conditioning set $C\subseteq\mathcal{V}$ if: (i) $a$ and $b$ are the endpoints; (ii) all non-colliders are not in $C$; and (iii) all colliders are in $\operatorname{an}_{G}(C)$. Nodes $a$ and $b$ are m-separated given $C$ if no such path exists.
@@
where $\text{sib}_{\mathcal{G}}(v)=\{u\in\mathcal{V}\mid\exists w\in\mathcal{V}:w\to v,\;w\to u,\;u\not\to v,\;v\not\to u\}$. We refer to [26] for multiple examples.
@@ 3.1 Connecting to Causality
In this perspective, we present the formulation of the Mixed Integer Quadratic Program (MIQP) used to infer the causal structure, with a new binary matrix $B=[b_{j,k}]\in\{0,1\}^{d\times d}$ introduced to account for relationships explained by confounding factors, alongside the weight matrix $W$ and binary adjacency matrix $E$ adopted from the ExDAG model [49]. Weight matrix $W$ represents the model weights. Whenever an entry in $W_{j,k}$ is non-zero, the respective value in either $E_{j,k}$ (directed edge) or $B_{j,k}$ is nonzero (bidirected edge). At the same time, we extend the existing formulation by introducing an additional binary input parameter $f_{j,k}$ for each pair of variables $(j,k)$, where $j\neq k$, indicating that there is no direct causal relationship between variables $j$ and $k$, but $j$ and $k$ might have a common cofactor. This follows from the meaning of the edges in a MAG $\rightarrow$ edge implies a direct causal relationship but does not rule out a possible latent confounding, $\leftrightarrow$ means no direct causal relationship.
Formally, we define Directed Edge Matrix $E$ as $e_{j,k}=1$ if $j\to k$, 00 otherwise. Similarly, Bidirected Edge Matrix $B$ is $b_{j,k}=1$, and 00 otherwise. Lastly, input matrix $F$ with the set of variables known to be independent, is by definition $f_{j,k}=1$ if $j\not\to k\wedge k\not\to j$ and 00 otherwise. Based on these assumptions, one can substitute (3) into (2), and the work of [49] is extended as explained in the next section.